# ChartOptimiser: Task-driven Optimisation of Chart Designs

### v1

@@ Abstract.
Effective chart design is essential for satisfying viewers’ information needs, such as retrieving values from a chart or comparing two values. However, creating effective charts is challenging and time-consuming due to the large design space and the inter-dependencies between individual design parameters. To address this challenge, we propose ChartOptimiser – a Bayesian approach for task-driven optimisation of charts, such as bar charts. At the core of ChartOptimiser is a novel objective function to automatically optimise an eight-dimensional design space combining four perceptual metrics: visual saliency, text legibility, colour preference, and white space ratio. Through empirical evaluation on 12 bar charts and four common analytical tasks – finding the extreme value, retrieving a value, comparing two values, and computing a derived value – we show that ChartOptimiser outperforms existing design baselines concerning task-solving ease, visual aesthetics, and chart clarity. We also discuss two practical applications of ChartOptimiser: generating charts for accessibility and content localisation. Taken together, ChartOptimiser opens up an exciting new research direction in automated chart design where charts are optimised for users’ information needs, preferences, and contexts.
@@
ccs: Human-centered computing Information visualizationccs: Human-centered computing HCI theory, concepts and models
Figure 1. ChartOptimiser is the first method to optimise parameters of chart designs for users’ visual analytical tasks. Taking a plain chart and an analytical task as input, ChartOptimiser uses Bayesian optimisation to automatically adapt the chart design with respect to a novel objective function that combines four perceptual metrics.
[…]
This work contributes ChartOptimiser – a task-driven optimisation pipeline of chart design, making charts tightly geared to analytical tasks.
@@ 1. Introduction
The ability to effectively communicate information through visualisations is crucial in the information age (Börner et al., 2019; Pozdniakov et al., 2023). For example, information visualisations on social media are designed to communicate the goals or intended messages of those creating them (Elzer et al., 2011). Well-designed charts not only satisfy viewers’ information needs by enabling, e.g., easy retrieval or comparison of data values, but they can also enhance overall comprehension and decision-making processes (Borkin et al., 2013, 2015; Wang et al., 2022). However, creating effective visualisations remains challenging and time-consuming – visualisation designers must navigate a vast design space in which each design choice impacts others.
@@
One challenge of chart design is the inter-dependencies between various design parameters, such as colour schemes, chart types, data scaling, or labels. While the choice of parameters is crucial for the overall effectiveness of charts, these choices can also lead to misinterpretations if they are not carefully made. The need to balance functional clarity with aesthetics further increases the difficulty of chart design. Visual appeal can only be evaluated by "taking a look at" the chart, i.e., viewers produce the image only when evaluating it. Consequently, designers must understand visual perception principles and data characteristics to create understandable, informative, and effective visualisations. Finally, it is challenging to design visualisations for the various analytical tasks that users may face when trying to satisfy their information needs (Sedlmair et al., 2011; Ziemkiewicz et al., 2020). Analytical tasks are among, if not the most important, factors influencing how viewers visually explore, make sense of, and extract information from visualisations.
The ever-increasing importance of information visualisations across various areas and the significant challenges in designing them call for automatic approaches that allow designers to optimise charts according to viewers’ needs. However, few previous works explored automatic optimisation of charts. Some works instead focused on automatically generating visualisations either using machine learning algorithms or user-defined rules and visual embellishments (Zhu et al., 2020). These methods have been studied in various application contexts ranging from extracting reusable layout templates from visualisations (Chen et al., 2020), over generating visualisations from natural language (Narechania et al., 2020; Luo et al., 2021), to optimising the layout according to various metrics (Micallef et al., 2017; Wu et al., 2021). However, no previous work has integrated viewers’ visual attention into this optimisation process, even though attention is well-known to convey rich information about analytical tasks and memory recall (Bulling et al., 2011; Bulling and Roggen, 2011; Wang et al., 2024b), as well as decision-making processes (Strauch et al., 2021; Kim et al., 2012).
@@
Optimising visualisations to satisfy users’ information needs is usually an iterative process that requires evaluating the visualisations at each iteration, which is inefficient and costly because the problem of finding the optimal parameter hypercube is inherently challenging, and the design space is vast. Recently, Bayesian optimisation (Shahriari et al., 2015) has shown good performance in optimising costly black-box problems and has been demonstrated to be highly effective for various applications such as optimising map searching (Dudley et al., 2019) and accelerating materials design (Zhang et al., 2020). Despite the effectiveness of Bayesian optimisation, it has not been explored for optimising visualisations.
@@
We propose ChartOptimiser – a novel Bayesian approach for task-driven optimisation of chart design that considers human visual attention. At the core of ChartOptimiser is a novel objective function that combines four perceptual metrics to evaluate and automatically identify the optimal chart design parameters within an eight-dimensional design space: highlighting saliency, text legibility, colour preference, and white space ratio. We demonstrate the usefulness of our method on the sample task of optimising the design parameters of bar charts – a widely used chart type in information visualisation research and practice. Through empirical evaluation of our method on 12 bar charts and four common visual analytical tasks – find extremum, retrieve value, compare, and compute derived value – we show that these four objectives are sufficient for effective optimisation, contrasting prior work that required many more objectives (Micallef et al., 2017). We further demonstrate that our ChartOptimiser outperforms existing design baselines concerning task-solving ease and clarity of the resulting chart. Our method also shows competitive performance with human-designed charts in terms of aesthetics and surpasses all baseline methods, particularly in more complex analytical tasks that involve comparison or computing derived values. In summary, the contributions of our work are two-fold:
@@
    We introduce ChartOptimiser  – the first automatic method to optimise chart design parameters to users’ analytical tasks. Our method uses Bayesian optimisation with a novel objective function that combines four perceptual metrics. We show that these metrics are sufficient for effective optimisation. We further discuss two sample applications that underline the practical usefulness of ChartOptimiser: designing charts for accessibility and content localisation.
@@
    We report an empirical comparison of ChartOptimiser with four baseline design methods that demonstrate our method’s superiority and the high similarity of the optimised visualisations to those created by human designers.
@@ 2.1. Perceptual Metrics of Information Visualisation
It is indubitable that the effective use of colours is a key factor in any visual design, including information visualisations (Rhyne, 2017; Szafir, 2017). Palmer and Schloss (2010) proposed the weighted affective valence estimates (WAVE) score to quantify the colour harmony and aesthetics in visualisation design. (Hegemann et al., 2023) proposed an integrated technique for colour exploration, assignment, and refinement, which took colour harmony, visual saliency, and elementary accessibility requirements into consideration. Visual saliency, defined as a measure of the regions in a visual stimulus or scene that attracts viewer attention, has been extensively studied, leading to the development of numerous effective saliency models in natural scenes (Itti et al., 1998; Harel et al., 2007; Kümmerer et al., 2017). However, saliency models have been found to perform poorly when applied to information visualisations, as highlighted in prior work (Matzen et al., 2017; Wang et al., 2024a). This triggered several research developing saliency models tailored for information visualisations in both free-viewing (Matzen et al., 2017; Shin et al., 2022; Wang et al., 2024a) and task-driven settings (Wang et al., 2024c). Some other perceptual metrics are specifically designed for information visualisations. Data-ink ratio quantifies the share of ink presenting data-related information in a visualisation (Tufte et al., 1990). Visual density rates the overall density of visual elements without distinguishing between data and non-data elements (Borkin et al., 2013). Sedlmair and Aupetit (2015) evaluated how separable different classes were in projections of high-dimensional datasets. Micallef et al. (2017) proposed three perceptual metrics devised for scatterplots, including the perception of linear correlation, image quality, and classes and outliers. As an integration of perceptual metrics, Shin et al. (2023) proposed the Virtual Human Visual System, which provides accessibility, text, visual saliency, and visual density as feedback for iterative visualisation design, enabling designers to refine their visualisations better to align saliency with their design intentions. Although the Virtual Human Visual System is promising, the involvement of visualisation designers is inevitable. Instead, this paper builds a fully automatic layout optimisation pipeline using four perceptual metrics in the objective function.
@@ 2.2. Automatic Visualisation Design
Automatic design systems could enumerate design candidates within the design space and find the best layout to satisfy the needs of designers or users. It usually requires a utility or objective function to rank the design candidates. Researchers have proposed many systems (Moritz et al., 2018; Hu et al., 2019; Luo et al., 2018) that recommend visualisations based on data structures and characteristics. Those systems focus on deciding the effective chart type, visual encoding, and data transformation. In addition to effectiveness, much research has optimised visualisations from other aspects. For example, Hopkins et al. (2020) addresses data integrity by surfacing chart construction errors such as truncated axes. Previous works have automatically extracted reusable layout templates from visualisations (Chen et al., 2020), sketches (Song et al., 2022a), and infographics (Lu et al., 2020). Other systems optimise the layout according to various metrics, such as similarities with user-input layouts (Tang et al., 2020), handcrafted energy terms, e.g. white space, scale, alignment, and balance (O’Donovan et al., 2014), perceptual metrics (Micallef et al., 2017), and crowdsourced layout quantifier (Wu et al., 2021). Generating visualisations from natural language (NL2Vis) has been a trend of automatic visualisation for many years (Narechania et al., 2020; Luo et al., 2021; Song et al., 2022b). NL2Vis takes a tabular dataset and a natural language query as input, specifying tasks and visualisation type, and generates visualisations that meet the requirement (Narechania et al., 2020; Luo et al., 2021). Recently, several works have demonstrated the superior performance of large language models (e.g. ChatGPT) for NL2Vis, highlighting their great potential under in-domain and cross-domain settings (Wu et al., 2024; Chen et al., 2024). However, current methods do not consider human perception feedback, such as visual attention. This work optimises chart design for a combination of four perceptual metrics, making the optimised charts closer to observation conditions and human perception.
@@ 2.3. Visualisation Analytical Tasks
Strong evidence has shown that tasks significantly influence how viewers interact with visualisations (Polatsek et al., 2018; Wang et al., 2024c). Amar et al. (2005) identified 10 low-level analytical tasks (e.g., retrieving value, finding extremum, comparison) while another study (Hibino, 1999) highlighted high-level tasks such as background understanding, planning of analysis, and data exploration. Saket et al. (2018) evaluated the effectiveness of basic types of visualisations in the 10 low-level analytical tasks. Elzer et al. (2011) proposed a Bayesian network to identify the communicative signals for intention recognition on simple bar charts. Eye tracking data collected by Polatsek et al. (2018) under three low-level analytical tasks demonstrated consistent gaze patterns within each task, with significant differences across tasks. Similarly, Wang et al. (2024c) found saliency metrics significantly varied across different tasks. Considering specific tasks is crucial, as visualisations can be designed to support a range of tasks based on the input data and can be evaluated by how effectively they enable task completion (Schulz et al., 2013). To support task execution in visualisations, some researchers have developed visualisations that explicitly highlight “data facts”, such as trends or comparisons (Srinivasan et al., 2018; Shi et al., 2020). In this work, we take task saliency as a metric in the objective function of the optimisation pipeline.
@@ 3. ChartOptimiser
ChartOptimiser aims to automate the optimisation of chart design parameters for a given analytical task (see Figure 2). Our method operates on a design space consisting of eight dimensions and optimises these with respect to a novel objective function that combines four perceptual metrics. Bayesian optimisation is used to sample from the design space and maximise the objective function. In the following, we describe each of these components in more detail. Our dataset and code are publicly available at https://xxx.xxxxxxxxxxx.xx.
Figure 2. Overview of ChartOptimiser. Our method operates on a design space with eight dimensions that it optimises with respect to an objective function consisting of four perceptual metrics. Bayesian optimisation is used to sample from the design space and to maximise the objective function, thus optimising the chart design for a given visual analytical task.
[…]
Overview of ChartOptimiser. Our method operates on a design space with eight dimensions that it optimises with respect to an objective function consisting of four perceptual metrics. Bayesian optimisation is used to sample from the design space and to maximise the objective function, thus optimising the chart design for a given visual analytical task.
@@ 3.1. Design Space
The design of charts involves two main components (Wu et al., 2021): 1) visual encodings that map data items to visual representations such as colour, position, and size, and 2) visual styles that specify data-irrelevant visual designs like orientation and bar width. This work focuses on chart design space for visual styles, specifically for bar charts. Bar charts are a widely used visualisation type, mainly for two-dimensional data. It divides the categorical data values into distinct bars, with values encoded by the bar length. Thus, categorical labels on axes associated with each bar are essential to a meaningful bar chart. Moreover, data labels on or next to bars showing the exact value of bars are also commonly used.
For task-driven chart design, highlighting the bar (s) relevant to tasks was shown to facilitate task completion (Elzer et al., 2011). For this initial work, we only include single-group bar charts to reduce the design space to a reasonable size. Key parameters in the design space of bar charts are (see Figure 2): 0) the aspect ratio of the chart, the font size of 1) axis labels and 2) data labels, 3) the bar width and 4) colour, often a 5) highlight bar colour, 6) the label rotation, and 7) the orientation of the chart. This set of design parameters extends the design space proposed in previous work (Wu et al., 2021).
We used the Vega-Lite (Satyanarayan et al., 2016) grammar for interactive graphics to specify the data visualisations. Vega-Lite specifications describe charts as encoding mappings from data to visual properties of marks, which usually requires a given mark type and a set of one or more encoding definitions for the visual channels. Vega-Lite expects a relational table of records with named fields. The mark type specifies the geometric objects used to encode data records visually. Possible values include bar, point, area, line, and tick. Visual encoding determines how data values map to the visual properties of marks. An encoding uses a visual channel such as spatial position, colour, size, shape, or text. The Vega-Lite compiler will infer defaults based on the channel and data type if omitted. With the Vega-Lite descriptive grammar, the entire design space of charts is represented as
@@
where $\mathcal{X}_{1}$ to $\mathcal{X}_{n}$ each denotes a dimension in the design space, such as aspect ratio and colour encoding. The $V$ function is the Vega-Lite compiler that automatically renders a chart from certain design parameters and determines the default properties for missing components based on a set of carefully designed rules (Satyanarayan et al., 2016). A certain parameter set $\mathbf{x}=(x_{1},x_{2},...x_{n}),x_{i}\in\mathcal{X}_{i}$ corresponds to one chart, denoted as $V_{\mathbf{x}}$.

### v2

@@ Abstract.
Automated chart design has seen significant advancements with the emergence of Large-Language Models (LLMs), which offer a practical solution for generating charts. However, LLMs frequently introduce possibly critical design failures, such as data manipulation and confabulation. While expert users can potentially mitigate these issues through iterative prompt engineering, this process requires substantial design knowledge and significant effort, remaining a massive barrier for the general public. In this paper, we present ChartOptimiser, an automated method for generating chart designs with fidelity, efficiency, and effectiveness. Given the inter-dependencies between individual design parameters, ChartOptimiser employs Bayesian optimisation to effectively search the chart design space for a novel objective function grounded in four perceptual metrics. Our empirical evaluations in bar and pie charts demonstrate that ChartOptimiser eliminates iterative design loops, providing non-expert users with high-quality charts that outperform LLM-generated designs in chart clarity, task-solving ease, and visual aesthetics.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title; xxxx xx–$x$, 2026; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Information visualizationccs: Human-centered computing HCI theory, concepts and models
Figure 1. While Large-Language Models demonstrate broad competence, their application to visualisation often results in failures – such as data manipulation, unsorted data, or low data-ink ratios – unless guided by expert prompting. In contrast, ChartOptimiser guarantees efficient, effective, and trustworthy designs without requiring design expertise from users. By taking a plain chart and an analytical task as input (e.g. “comparing CO2 emissions”), ChartOptimiser employs Bayesian optimisation to effectively maximise a novel objective function composed of four perceptual metrics on visualisations.
This work contributes ChartOptimiser that employs Bayesian optimisation to effectively maximise a novel objective function composed of four perceptual metrics.
@@ 1. Introduction
The ability to effectively communicate information through visualisations is crucial in the information age (Börner et al., 2019; Pozdniakov et al., 2023). Well-designed charts not only satisfy viewers’ information needs by enabling, e.g. easy retrieval or comparison of data values, but they can also enhance overall comprehension and decision-making processes (Borkin et al., 2013, 2015; Wang et al., 2022). However, creating effective visualisations remains challenging and time-consuming – visualisation designers must navigate a vast design space in which each design choice impacts others.
@@
One challenge of chart design is the inter-dependencies between various design parameters, such as colour schemes, chart types, data scaling, or labels. While the choice of parameters is crucial for the overall effectiveness of charts, these choices can also lead to misinterpretations if they are not made carefully. The need to balance functional clarity with aesthetics further increases the difficulty of chart design. Visual appeal can only be evaluated by “taking a look at” the chart; that is, viewers produce the image only when evaluating it. Consequently, designers must understand visual perception principles and data characteristics to create understandable, informative, and effective visualisations. Finally, it is challenging to design visualisations for the various analytical tasks that users may face when trying to satisfy their information needs (Sedlmair et al., 2011; Ziemkiewicz et al., 2020). Analytical tasks are among the most important factors influencing how viewers visually explore, make sense of, and extract information from visualisations.
The ever-increasing importance of data visualisations across various areas and the significant challenges in designing them call for automatic approaches that allow designers to optimise charts according to viewers’ needs. Some works focused on automatically generating visualisations either using machine learning algorithms (Chen et al., 2019; Zhao et al., 2020) or rule-based constrains (Wang et al., 2024a). These methods have been studied in various application contexts ranging from extracting reusable layout templates from visualisations (Chen et al., 2020), over generating visualisations from natural language (Narechania et al., 2020; Luo et al., 2021), to optimising the layout according to various metrics (Micallef et al., 2017; Wu et al., 2021). Recently, automated chart design has seen significant advancements with the emergence of Large-Language Models (LLMs), which offer a seamless solution for generating charts (Tian et al., 2024; Wang et al., 2024a). However, confabulation (Siontis et al., 2024)(also known as hallucination) is a longstanding problem that LLMs manipulates the data and compromises trustworthiness. Compared with the design constraints in rule-based chart optimisation methods, LLMs generate charts from billions of tokens that updated from “in-the-wild” data sources, leading to charts that may look plausible but violate chart design principles. While expert users can potentially mitigate these issues through iterative prompt engineering, this process requires substantial design knowledge and significant effort that remains a massive barrier for the general public.
To address these limitations, we propose a shift towards Bayesian optimisation to generate chart designs with the following advantages:
  • Fidelity: Bayesian optimisation searches within a constrained parameter space of valid chart specifications (e.g. VegaLite), guaranteeing data integrity and preventing data manipulation.
  • Efficiency: Bayesian optimisation reduces the manual effort of the iterative design revisions, such as prompt engineering to improve or correct LLM outputs.
  • Effectiveness: While LLMs often focus on aesthetic plausibility, Bayesian optimisation directly maximises a novel objective function grounded in perceptual metrics, ensuring that the resulting visualisations are empirically effective for specific analytical tasks.
@@
Optimising visualisations to satisfy users’ information needs is usually an iterative process that requires evaluating the visualisations at each iteration, which is inefficient and costly because the problem of finding the optimal parameter hypercube is inherently challenging, and the design space is vast. Recently, Bayesian optimisation (Shahriari et al., 2015) has shown good performance in optimising costly black-box problems and has been demonstrated to be highly effective for various applications such as optimising map searching (Dudley et al., 2019) and accelerating materials design (Zhang et al., 2020). However, it has not yet been explored for optimising chart designs to specific analytical tasks.
@@
We propose ChartOptimiser – a novel Bayesian approach for task-driven optimisation of chart design that considers human visual attention. At the core of ChartOptimiser is a novel objective function that combines four perceptual metrics to evaluate and automatically identify the optimal chart design parameters within an eight-dimensional design space: highlighting saliency, text legibility, colour preference, and white space ratio. We demonstrate the usefulness of our method on the sample task of optimising the design parameters of bar charts – a widely used chart type in data visualisation research and practice. Through empirical evaluation of our method on 12 bar charts and four common visual analytical tasks – find extremum, retrieve value, compare, and compute derived value – we show that these four objectives are sufficient for effective optimisation, contrasting prior work that required many more objectives (Micallef et al., 2017). We further demonstrate that our ChartOptimiser outperforms existing design baselines concerning task-solving ease and clarity of the resulting chart. Our method also shows competitive performance with human-designed charts in terms of aesthetics and surpasses baseline methods, particularly in more complex analytical tasks that involve comparison or computing derived values. In summary, the contributions of our work are two-fold:
@@
    We introduce ChartOptimiser  – a fully automatic method to optimise chart design parameters to users’ analytical tasks. Our method uses Bayesian optimisation with a novel objective function that combines four perceptual metrics. We further showcased how ChartOptimiser could be adapted to many chart types such as multi-column bars and pie charts.
@@
    We report an empirical comparison of ChartOptimiser with four baseline design methods, including large-language models (e.g. ChatGPT), demonstrating our method’s ability of fidelity, efficiency, and effectiveness in chart design optimisation.
@@ 2.1. Perceptual Metrics of Information Visualisation
It is indubitable that the effective use of colours is a key factor in any visual design, including information visualisations (Rhyne, 2017; Szafir, 2017). Palmer and Schloss (2010) proposed the weighted affective valence estimates (WAVE) score to quantify the colour harmony and aesthetics in visualisation design. (Hegemann et al., 2023) proposed an integrated technique for colour exploration, assignment, and refinement, which took colour harmony, visual saliency, and elementary accessibility requirements into consideration. Visual saliency, defined as a measure of the regions in a visual stimulus or scene that attract viewer attention, has been extensively studied, leading to the development of numerous effective saliency models in natural scenes (Itti et al., 1998; Harel et al., 2007; Kümmerer et al., 2017). However, saliency models have been found to perform poorly when applied to information visualisations, as highlighted in prior work (Matzen et al., 2017; Wang et al., 2024b). This triggered several research developing saliency models tailored for information visualisations in both free-viewing (Matzen et al., 2017; Shin et al., 2022; Wang et al., 2024b) and task-driven settings (Wang et al., 2024d). Some other perceptual metrics are specifically designed for information visualisations. Data-ink ratio quantifies the share of ink presenting data-related information in a visualisation (Tufte et al., 1990). Visual density rates the overall density of visual elements without distinguishing between data and non-data elements (Borkin et al., 2013). Sedlmair and Aupetit (2015) evaluated how separable different classes were in projections of high-dimensional datasets. Recallability and recognisability quantifies the quality of infographics design by testing how much information viewers remember after observation (Borkin et al., 2015; Wang et al., 2022, 2024c). Micallef et al. (2017) proposed three perceptual metrics devised for scatterplots, including the perception of linear correlation, image quality, and classes and outliers. As an integration of perceptual metrics, Shin et al. (2023) proposed the Virtual Human Visual System, which provides accessibility, text, visual saliency, and visual density as feedback for iterative visualisation design, enabling designers to refine their visualisations better to align saliency with their design intentions. Although the Virtual Human Visual System is promising, the involvement of visualisation designers is inevitable. Instead, this paper builds a fully automatic design optimisation pipeline that incorporates four perceptual metrics into the objective function, without any human effort in the design loop.
@@ 2.2. Automatic Visualisation Design
Automatic design systems can enumerate design candidates within the design space and identify the optimal layout to meet the needs of designers or users. It usually requires a utility or objective function to rank the design candidates. Researchers have proposed many systems (Moritz et al., 2018; Hu et al., 2019; Luo et al., 2018; Zhao et al., 2020) that recommend visualisations based on data structures and characteristics. Those systems focus on deciding the effective chart type, visual encoding, and data transformation. In addition to effectiveness, much research has optimised visualisations from other aspects. For example, Hopkins et al. (2020) addresses data integrity by surfacing chart construction errors such as truncated axes. Previous works have automatically extracted reusable layout templates from visualisations (Chen et al., 2020), sketches (Song et al., 2022a), and infographics (Lu et al., 2020). Other systems optimise the layout according to various metrics, such as similarities with user-input layouts (Tang et al., 2020), handcrafted energy terms, e.g. white space, scale, alignment, and balance (O’Donovan et al., 2014), perceptual metrics (Micallef et al., 2017), and crowdsourced layout quantifier (Wu et al., 2021). Generating visualisations from natural language (NL2Vis) has been a trend in automatic visualisation for many years (Narechania et al., 2020; Luo et al., 2021; Song et al., 2022b; Guo et al., 2024). NL2Vis takes a tabular dataset and a natural language query as input, specifying tasks and visualisation type, and generates visualisations that meet the requirement (Narechania et al., 2020; Luo et al., 2021; Cui et al., 2019). Recently, several works have demonstrated the superior performance of large language models (e.g. ChatGPT) for NL2Vis, highlighting their great potential in both in-domain and cross-domain settings (Wu et al., 2024; Chen et al., 2024). However, current methods do not consider human perception feedback, such as visual attention. This work optimises chart design for a combination of four perceptual metrics, making the optimised charts closer to observation conditions and human perception.
@@ 2.3. Visualisation Analytical Tasks
Strong evidence has shown that tasks significantly influence how viewers interact with visualisations (Polatsek et al., 2018; Wang et al., 2024d). Amar et al. (2005) identified 10 low-level analytical tasks (e.g., retrieving value, finding extremum, comparison) while another study (Hibino, 1999) highlighted high-level tasks such as background understanding, planning of analysis, and data exploration. Saket et al. (2018) evaluated the effectiveness of basic types of visualisations in the 10 low-level analytical tasks. Elzer et al. (2011) proposed a Bayesian network to identify the communicative signals for intention recognition on simple bar charts. Eye movement data by Polatsek et al. (2018) and model by Shi et al. (2025) under three low-level analytical tasks demonstrated consistent gaze patterns within each task, with significant differences across tasks. Similarly, Wang et al. (2024d) found saliency metrics are important indicators for visual analytical tasks. Considering specific tasks is crucial, as visualisations can be designed to support a range of tasks based on the input data and can be evaluated by how effectively they enable task completion (Schulz et al., 2013). To support task execution in visualisations, researchers have developed visualisations that explicitly highlight “data facts”, such as trends or comparisons (Srinivasan et al., 2018; Shi et al., 2020, 2021, 2024). In this work, we take task saliency as a metric in the objective function of the optimisation pipeline.
@@ 3. ChartOptimiser
ChartOptimiser aims to automate the optimisation of chart design parameters for a given analytical task (see Figure 2). Our method operates on a design space consisting of eight dimensions and optimises these with respect to a novel objective function that combines four perceptual metrics. Bayesian optimisation is used to sample from the design space and maximise the objective function. In the following, we describe each of these components in more detail. Our code is publicly available at https://anonymous.4open.science/r/ChartOpt-96AB.
Figure 2. Overview of ChartOptimiser. Our method operates in a design space with eight dimensions, optimising it with respect to an objective function consisting of four perceptual metrics. Bayesian optimisation is used to sample from the design space and to maximise the objective function, thus optimising the chart design for a given visual analytical task.
Overview of ChartOptimiser. Our method operates in a design space with eight dimensions, optimising it with respect to an objective function consisting of four perceptual metrics. Bayesian optimisation is used to sample from the design space and to maximise the objective function, thus optimising the chart design for a given visual analytical task.
@@ 3.1. Design Space
The chart design space involves two main components (Wu et al., 2021): 1) visual encodings that map data items to visual representations such as colour, position, and size, and 2) visual styles that specify data-irrelevant visual designs like orientation and bar width. This work mainly focuses on bar charts, as they are the most widely used chart type for two-dimensional data. It divides the categorical data values into distinct bars, with values encoded by the bar length. Thus, categorical labels on axes associated with each bar are essential to a meaningful bar chart. Moreover, data labels on or next to bars showing the exact value of bars are also commonly used.
For task-driven chart design, highlighting the bar (s) relevant to tasks was shown to facilitate task completion (Elzer et al., 2011). For this initial work, we have limited ourselves to single-group bar charts to reduce the design space to a manageable scale. Key parameters in the design space of bar charts are (see Figure 2): 0) the aspect ratio of the chart, the font size of 1) axis labels and 2) data labels, 3) the bar width and 4) colour, often a 5) highlight bar colour, 6) the label rotation, and 7) the orientation of the chart. This set of design parameters extends the design space proposed in previous work (Wu et al., 2021).
We used the VegaLite (Satyanarayan et al., 2016) grammar for interactive graphics to specify the data visualisations. VegaLite specifications describe charts as encoding mappings from data to the visual properties of marks, which typically require a specific mark type and a set of one or more encoding definitions for the visual channels. VegaLite expects a relational table of records with named fields. The mark type specifies the geometric objects used to encode data records visually. Possible values include bar, point, area, line, and tick. Visual encoding determines how data values map to the visual properties of marks. An encoding uses a visual channel such as spatial position, colour, size, shape, or text. The VegaLite compiler will infer defaults based on the channel and data type if omitted. With the VegaLite descriptive grammar, the entire design space of charts is represented as
@@
where $\mathcal{X}_{1}$ to $\mathcal{X}_{n}$ each denotes a dimension in the design space, such as aspect ratio and colour encoding. The $V$ function is the VegaLite compiler that automatically renders a chart from certain design parameters and determines the default properties for missing components based on a set of carefully designed rules (Satyanarayan et al., 2016). A certain parameter set $\mathbf{x}=(x_{1},x_{2},...x_{n}),x_{i}\in\mathcal{X}_{i}$ corresponds to one chart, denoted as $V_{\mathbf{x}}$.